# Experimental demonstration of a robust training method for strongly defective neuromorphic hardware

### v1

@@ I Introduction
In this work, we study training methods on non-ideal arrays by fabricating 36 dies each containing 20,000 MTJs integrated with CMOS into a crossbar array structure, where each die represents the weights of a two-layer BNN. These arrays are described in Sec. II.1 and the network structure is described in Sec. II.2. Section III.1 reports that ex-situ training of each die where the unique defect characteristics of each die are considered recovers mean classification accuracy to within 2 % of the software baseline.
@@ II.1 Design and test of 20,000 MTJs integrated with CMOS
We design and fabricate a 2-transistor-1-resistor (2T-1R) test vehicle capable of housing 20,000 two-terminal devices in a crossbar array architecture shown in Figs. 1(a,b). The entire chip consists of 100 rows $\times$ 200 columns which are partitioned into four subarrays of 50 columns each. A two-bit digital signal selects the subarray, allowing for characterization of 5000 devices at a time. The two-transistor design acts as a pass gate circuit with an n-channel metal-oxide semiconductor (NMOS) and p-channel metal-oxide semiconductor (PMOS) transistor in parallel; a digital enable signal determines whether the pass-gate operates in a digital or analog mode. For all experiments in this work, measurements were performed in digital mode: the gate voltage on the NMOS (PMOS) transistor is power (ground), reducing as much as possible any series resistance with the MTJ. Initial designs for the 180-nm CMOS are fabricated in a commercial foundry without the MTJs and received with the final metal layer removed, exposing the tungsten (W) vias underneath.
@@ II.2 Neural network architecture design and mapping onto MTJ hardware
Under the current experimental setup we are not able to implement VMM in parallel on the network, which would require 100 unique voltages sourced from 100 unique source-measure unit channels. Work in Ref. [34] shows that if the devices are operated in a linear regime, a full VMM can be represented by applying a read voltage element-by-element to each device and summing the resulting currents in software. We have verified that the applied voltages for this work are in this regime and our measurements satisfy the superposition principle of linear circuits. For this reason, the following results are labeled as hardware emulation, because the conductances of MTJs are used in software to determine the output currents. Furthermore, due to the sheer volume of programming steps for 100 solutions across 36 dies, we take advantage of the very low cycle-to-cycle variation in MTJ resistance (see Appendix C) and program solutions onto the weight matrices in software using measured $R_{\text{P}}$ and $R_{\text{AP}}$ values obtained during screening of the dies.
@@ IV Summary and Outlook
In this work, 36 dies each containing a crossbar array of 20,000 CMOS-integrated MTJs are used to characterize the performance of a hardware implementation of a binary neural network on the MNIST dataset. We investigate the effect of strongly defective devices on the hardware implementation’s classification error in a context where large variation in performance between dies is observed. By implementing a hardware-aware training method that compensates for defective device locations by strengthening the weights in other information pathways, we show that the error for each die can recover to levels comparable to that of an ideal, defect-less crossbar.
@@ V ACKNOWLEDGEMENTS
WAB and AM contributed equally to this work. We thank Thomas Boone for fruitful discussions. This work was funded by the National Institute of Standards and Technology. Advait Madhavan acknowledges support under the Cooperative Research Agreement Award No. 70NANB14H209 through the University of Maryland and NSF grant number CCF-CISE-ANR-FET-2121957.
@@ Appendix A Device-to-device variation impact on classification error
We investigate the effect that device-to-device variation has on the classification error of the hardware emulation. In the main text, it is stated that the discrepancy between the performance of the hardware emulation and the defect-free solution for the hardware-aware training method is due to device-to-device variations inducing a distribution around the ternary weights. To show this, we simulate the MTJ substrate having various levels of variation in resistance. For each simulation, a crossbar is generated by sampling normal distributions with mean equal to the mean P and AP resistance states of the experimentally measured hardware and standard deviation determined by the coefficient of variation, defined as $\sigma=\alpha\times\mu$, where $\sigma$ is the standard deviation, $\alpha$ is the coefficient of variation, and $\mu$ is the mean. This generated crossbar is then used in the same manner as the other results presented in this paper. Fig. 6 shows the classification error of the test dataset for 100 pure software solutions (without defects) each on a crossbar with increasing levels of variation. As expected, at 0 % variation, the error does not differ from the software performance since there are only two values that the MTJs can represent. However as the variation increases, the discrepancy with the defect-free performance begins to diverge. It is worth noting that at 8 % variation, the discrepancy of the simulated hardware with defect-free performance is of the same level as that of the hardware-aware results shown in Fig. 2(e).
@@ Appendix C Yield statistics of all dies and cycle-to-cycle variation of MTJ resistance
In this appendix, we first show cycle-to-cycle variability of the resistance of each MTJ within a single die followed by a plot of heat maps showing the device yield of each die. As mentioned in the main text, inference is performed by emulating the 36 experimentally measured dies. Performing inference on each solution uses identical values for resistance, where in a completely hardware implementation, resistance values will vary. To investigate the levels of resistance variation of the dies in this work, we perform 100 cycles where each MTJ is passed through the screening test mentioned in II.1. Fig. 8 plots the coefficient of variation (defined as the standard deviation divided by the mean) for the P-state resistance $\alpha_{R}$ plotted for each of the 20,000 MTJs. This cycle-to-cycle variation test is performed for Die 1 and Die 36, radially on opposite ends of the wafer, to show the uniformity of variation. While only P-state resistance variation is shown here, AP-state resistance shows similar levels of variation. In 100 cycles, the resistance of the MTJ varies less than 5 %, suggesting this variation has a trivial contribution to any variation in performance of inference.

### v2

# Measurement-driven neural-network training for integrated magnetic tunnel junction arrays

@@ I Introduction
In this work, we study offline training methods on non-ideal arrays by fabricating 36 dies each containing 20,000 MTJs integrated with CMOS into a crossbar array structure, where each die represents the weights of a two-layer BNN. These arrays are described in Sec. II.1 and the network structure is described in Sec. II.2. Section III.1 reports that ex-situ training of each die where the unique defect characteristics of each die are considered recovers mean classification accuracy to within 2 % of the software baseline.
@@ II.1 Design and test of 20,000 MTJs integrated with CMOS
We design and fabricate a 2-transistor-1-resistor (2T-1R) test vehicle capable of housing 20,000 two-terminal devices in a crossbar array architecture shown in Figs. 1(a,b). The entire chip consists of 100 rows $\times$ 200 columns which are partitioned into four subarrays of 50 columns each. A two-bit digital signal selects the subarray, allowing for characterization of 5000 devices at a time. The two-transistor design acts as a pass gate circuit with an n-channel metal-oxide semiconductor (NMOS) and p-channel metal-oxide semiconductor (PMOS) transistor in parallel; a digital enable signal determines whether the pass-gate operates in a digital or analog mode.
The choice of a 2T-1R structure as opposed to a 1T-1R or passive array structure was made to produce a general test vehicle for characterizing many two-terminal novel device technologies. When the device properties call for current compliance, the PMOS can be turned off and the NMOS can be used. For all experiments in this work, measurements were performed in digital mode: the gate voltage on the NMOS (PMOS) transistor is power (ground), reducing as much as possible any series resistance with the MTJ. For ideal MTJs with narrow property distributions, a passive array is the best option for the highest density on chip. In a practical situation where device properties are well characterized but non-idealities such as sneak paths and line resistance exist, a final product of this chip would use a 1T-1R structure to isolate sections of the chip for measurements. The 2T-1R initial designs for the 180-nm CMOS are fabricated in a commercial foundry without the MTJs and received with the final metal layer removed, exposing the tungsten (W) vias underneath.
@@ II.2 Neural network architecture design and mapping onto MTJ hardware
Under the current experimental setup, we are not able to implement VMM in parallel on the network, which would require 100 unique voltages sourced from 100 unique source-measure unit channels. Work in Ref. [34] shows that if the devices are operated in a linear regime, a full VMM can be represented by applying a read voltage element-by-element to each device and summing the resulting currents in software. We have verified that the applied voltages for this work are in this regime and our measurements satisfy the superposition principle of linear circuits. For this reason, the following results are labeled as hardware emulation, because the conductances of MTJs are used in software to determine the output currents. Furthermore, due to the sheer volume of programming steps for 100 solutions across 36 dies, we take advantage of the very low cycle-to-cycle variation in MTJ resistance (see Appendix C) and program solutions onto the weight matrices in software using measured $R_{\text{P}}$ and $R_{\text{AP}}$ values obtained during screening of the dies.
@@ IV Summary and Outlook
In this work, 36 dies each containing a crossbar array of 20,000 CMOS-integrated MTJs are used to characterize the performance of a hardware emulation of a binary neural network on the MNIST dataset. We investigate the effect of strongly defective devices on the hardware’s classification error in a context where large variation in performance between dies is observed. By implementing a hardware-aware training method that compensates for defective device locations by strengthening the weights in other information pathways, we show that the error for each die can recover to levels comparable to that of an ideal, defect-less crossbar.
@@ V ACKNOWLEDGEMENTS
WAB and AM contributed equally to this work. We thank Thomas Boone for fruitful discussions. This work was funded by the National Institute of Standards and Technology. AM acknowledges support under the Cooperative Research Agreement Award No. 70NANB14H209 through the University of Maryland and NSF grant number CCF-CISE-ANR-FET-2121957.
@@ Appendix A Device-to-device variation impact on classification error
We investigate the effect that device-to-device variation has on the classification error of the hardware emulation. In the main text, it is stated that the discrepancy between the performance of the hardware emulation and the defect-free solution for the hardware-aware training method is due to device-to-device variations inducing a distribution around the ternary weights. To show this, we simulate the MTJ substrate having various levels of variation in resistance. For each simulation, a crossbar is generated by sampling normal distributions with mean equal to the mean P and AP resistance states of the experimentally measured hardware and standard deviation determined by the coefficient of variation, defined as $\sigma=\alpha\times\mu$, where $\sigma$ is the standard deviation, $\alpha$ is the coefficient of variation, and $\mu$ is the mean. This generated crossbar is then used in the same manner as the other results presented in this paper.
@@
Fig. 6 shows the classification error of the test dataset for 100 pure software solutions (without defects) each on a crossbar with increasing levels of variation. As expected, at 0 % variation, the error does not differ from the software performance since there are only two values that the MTJs can represent. However as the variation increases, the discrepancy with the defect-free performance begins to diverge. It is worth noting that at 8 % variation, the discrepancy of the simulated hardware with defect-free performance is of the same level as that of the hardware-aware results shown in Fig. 2(e).
@@ Appendix C Yield statistics of all dies and cycle-to-cycle variation of MTJ resistance
In this appendix, we first show cycle-to-cycle variability of the resistance of each MTJ within a single die followed by a plot of heat maps showing the device yield of each die. As mentioned in the main text, inference is performed by emulating the 36 experimentally measured dies. Performing inference on each solution uses identical values for resistance, where in a completely hardware implementation, resistance values will vary. To investigate the levels of resistance variation of the dies in this work, we perform 100 cycles where each MTJ is passed through the screening test mentioned in II.1.
Fig. 8 plots the coefficient of variation (defined as the standard deviation divided by the mean) for the P-state resistance $\alpha_{R}$ plotted for each of the 20,000 MTJs. This cycle-to-cycle variation test is performed for Die 1 and Die 36, radially on opposite ends of the wafer, to show the uniformity of variation. While only P-state resistance variation is shown here, AP-state resistance shows similar levels of variation. In 100 cycles, the resistance of the MTJ varies less than 5 %, suggesting this variation has a trivial contribution to any variation in performance of inference.